OpenStudy (anonymous):

Help. Solve (3/x^2 + 3x) + (x+2/x+3) < (1/x) Show me how to figure it out. :)

OpenStudy (anonymous):

subtract \(\frac{1}{x}\) from both sides add the terms then check to see where the resulting fraction is negative

OpenStudy (anonymous):

\[\frac{3}{x^2+3x}+\frac{x+2}{x+3}-\frac{1}{x}<0\] \[\frac{x+1}{x+3}<0\] etc

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw